A car has a speed VA 4 m/s at point A, is moving with a constant acceleration a = 2 m/s reaches point B with VB 12 m/s. a) Calculate the time needed for the car to reach point B. b) Calculate the distance AB. c) The particle reaches point B with VB 12m/s ,moves with this speed, to reach C after 20 seconds. 1) What is with justification the nature of motion between B and C. 2) Calculate the distance BC d) At C, the driver applies a constant braking force ,in a uniformly decelerated inotion in order to stop the car at point D after 5, seconds. D Calculate the acceleration between C and D. 2) Deduce the distance CD. e) Calculate the distance AD and deduce the average speed between A and D.


1
Expert's answer
2021-05-28T07:16:44-0400

a)


"t=\frac{12-4}{2}=4\ s"

b)


"12^2-4^2=2(2)s\\s=32\ m"

c)


"s'=12(20)=240\ m"

d)


"a'=-\frac{12}{5}=-2.4\frac{m}{s^2}"

e)


"d=0.5(2.4)(5)^2=30\ m"

f)


"D=32+240+30=302\ m"

g)


"V=\frac{302}{4+20+5}=10.4\frac{m}{s}"